# Distributed Multisensor ISAC

### v1

@@ V-E Sensing Cycles – OFDMA Resource Allocation, Time-frequency Precoding and Link Adaptation
It seems beneficial to adjust ICAS radio access parameters and resources to maximize location and tracking parameters according to the current sensing mode. However, at the beginning of a reconnaissance mission, we generally have no a priori knowledge of the existence, position and speed of the target. We would need dedicated radar sensing cycles following gradual levels of perception. In a first surveillance and acquisition cycle, the environment is scanned and when a target is detected, quick, rough and sometimes even incorrect estimates of the target parameters is made (e.g., many false alarms). Then, in a second step, the ISAC radio parameters will be adjusted to provide more accurate and less ambiguous estimates. This can be achieved, e.g., by a higher sensing symbol repetition rate, a longer coherent processing interval (c.f. radar integration time) which goes along with higher Doppler resolution, and by beamforming which adapts to the directivity of propagation. However, since the dynamics and contributions of the multiple radio links depend on their relative positions with respect to the target, which change over time, ISAC-specific CSI estimation and feedback are needed for dynamic transmitter resource allocation, precoding, and link adaptation to support dynamic target tracking. Since this is a two-stage process based on previous CSI measurements, the effectiveness depends on the CSI pilot rate and the speed at which the channel changes.
@@
Although the next generation mobile radio may include dedicated tools for ISAC quality of service control, it seems promising to explore how existing mechanisms and tools for CSI estimation and signaling can be used for ISAC performance control. Traditionally, intersymbol interference in mobile communications caused by multipath is compensated at the receiver by an equalizer. Later, it was realized that precoding at the transmitter side maybe more effective, especially if multiple users are included. Over the decades of its maturation, mobile communications has developed sophisticated methods for precoding, and resource allocation. Link adaption has evolved into an unprecedentedly flexible multi-domain technique that includes scalable numerology, flexible resource element allocation in frequency and time, waveform selection, carrier power loading, modulation and coding, and multiple antenna precoding techniques including beamforming [ref47]. Together with advanced techniques for CSI estimation and signaling, the control channel allows to reconfigure and adapt the transmitter to the instantaneous link conditions for optimum quality of communication service. The latest stage has been set with 5G NR, which is designed to support enhanced mobile broadband, ultra-reliable low latency communications and massive machine type communications. Since coordination of multiple links and multipath exploitation will be even more important for performance optimization in ISAC (due to the geometric relevance of the multiple distributed links), we expect that precoding and link adaptation will be of similar or even greater importance.
@@
Let’s take another quick look at the link adaptation in Radar. The situation is slightly different here. Although multistatic and networked Radar have a long lasting history [ref49], single station monostatic radars are still the default design [ref48]. The reasons for this seem to be manifold. Monostatic radars are independent and can act autonomously (even mobile). The underlying geometry, a circle or a sphere, seems to be simple. Most of all, signal processing is local, which makes Tx/Rx synchronization and resource adaption (bandwidth, Tx power, pulse rate, dwell time, beam management etc.) a local task. There are very advanced resource management principles for monostatic radars applied that can cope with dynamic scenarios of multiple targets [ref46] and are able to adapt their operation based on cognitive procedures [ref44]. Attempts to overcome the geometric limitations of coverage and target visibility by a network of radar sensors often ends up with non-coherent cooperation of multiple monostatic radars [ref45] (fusion of multiple directions from dislocated radars is a non-coherent operation). References to radar networks with multi-static cooperation are rare [ref48]. Even if the publicly available literature may not completely reflect the entire state of knowledge, the radar-centric approach to ISAC, which requires a separate (probably proprietary) communication network or additional communication functions integrated into the radar’s air interface, does not appear to be a very promising role model for large scale application of ISAC in public networks. At the very least, ISAC’s communications-centric approach seems more promising, as it reuses the already available communication network for data transport and fusion. In addition, the well-designed and constantly evolving sensing, signaling, and control schemes of mobile communications reveal a kind of hitchhiking support for radar functionality. This is most obvious for the CPCL approach, which seems to be the closest coexistence of communication and sensing.
@@
As we have explained above, ISAC radar detection and tracking of 3D dynamic target state vector in general requires multiple sensing links that have to be coordinated and adapted to the dynamic target scenario to meet the required sensing performance. It is obvious that resource management and link adaptation must take place in a distributed MS ISAC system at the radio and network level. While the primary goal of communication is efficient transfer of data between two dedicated radio nodes acting as the source and the destination, the goal of ISAC is the localization and recognition of target objects. Multiuser communication links, in essence, can be seen as independent single links just sharing common radio resources according to rules of fairness or priority. Sensing rather requires the proactive cooperation of multiple radio nodes (illuminators and sensors) that share the same resources and work together to fulfill a common sensing task. Already when selecting the nodes involved, their spatial availability (both current and predicted), expected influence on localization accuracy due to GDoP and SNR, and detection functionality due to heterogeneous configuration and connectivity have to be taken into account. Furthermore, the allocation of communication and computation resources is based on the decision on the data collection task and the type of data fusion. It seems important to recognize that not a single radio node (e.g. a UE) is a beneficiary of an ISAC service, but the fused information from several sensing links that is made available to authorized users represents the added value. Therefore, MS ISAC resource allocation is always a joint, respectively cooperative activity of a sensing network.
@@
Once we have defined which radio nodes contribute to the execution of a sensing mission, we need to decide how to allocate radio access resources. This is especially interesting when multiple sensing links share the same radio channel. Which signals and access schemes do we have available for target scene illumination in ISAC? While dedicated pilot signals are required to establish communication links (Tx/Rx synchronization, demodulation reference signaling and CSI estimation), and therefore are essential also for ISAC operation, the use of CSI pilots alone will in general not be enough for ISAC sensing as already explained in the subsection on Recovery of the Correlation Reference Signal. The use of the Positioning Reference Signal would be another option. The PRS is a dedicated signal that enables self-localization of mobile UEs in the DL. It can carry network specific data to support localization services and allows OFDMA frame mapping to support multiple UEs [ref74]. Although it is not intended or radar sensing, it can be used in the DL of MS ISAC to temporarily enhance target illumination. However, excessive use of PRS would only block resources that may not be available for communication services. This means that radio resource efficient ISAC should aim to reuse the communication payload for sensing, as already proposed by the CPCL principle. It is also possible to use the communication payload based sensing results for CSI feedback and sensing link adaption. The use of CSI pilots has the advantage that orthogonal multi-sensor access is easy to achieve.
@@
According to the subsection on Multi-Sensor Radio Access and Scheduling it is reasonable to distinguish between the broadcast mode and multi-sensor ISAC access. In the first case, each transmitter serves several sensing receivers, which provides simultaneous measurements and, thus, instantaneously multiplies the information about the target. This applies, e.g., if the sensing nodes are receive-only sniffers deployed as part of a distributed infrastructure or acting as deployable sniffers in the reduced CPCL-mode as described in section II. In case of MS ISAC, the orthogonal multiplexed sensing links must share the same channel by proper allocation of the RBs in the OFDM frequency-time frame as described in the subsection on Multi-Sensor Radio Access and Scheduling. Obviously, resource sharing should be made adaptive according to the number and dynamics of targets, as well as the multipath propagation situation.
@@
For the purpose of MS ISAC radio access in DL or UL, we must adjust the distribution of RB in the frequency-time frame by OFDMA. From the instantaneous ISAC CSI and KPI estimates transmitted via the UL to the central unit (CU) of a distributed base station, a resource scheduler can decide how much resources per link have to be spent to follow the dynamically changing parameters of the respective target. Later on, when the moving target is tracked and the dynamic situation changes according to the targets position and velocity vector relative to the baseline of the sensor links, the RB distribution will be updated.
@@
Since the target resolution performance depends on the Cramer Rao Bound (CRB) as discussed in the subsection on Model based Bistatic Target Parameter Estimation, it recommends to use it as cost function for MS ISAC resource scheduling. For the MS access, the bottom line is about optimum arrangement and average power weighting of carriers in the OFDMA frame for any sensing link. The total number of carriers to be allocated per link depends on both the required processing gain of the matched filter (according to the SNR of the respective link) and the resolution performance. The CRLB criterion tends to shift the carriers in blocks to the edges of the available frequency-time frame. Since we may have to host multiple sensing links, we would need to interleave them in two dimensions, frequency and time. A simple role of thumb for RB scheduling of multiple cooperating sensing links may be to allocate the RBs of links with temporarily lower resolution requirements to the inner space of the OFDMA frame. The advantage of the 2D nesting is that virtual simultaneous measurements from multiple transmitters become possible - with the restrictions that a shared medium entails.
@@
The weighting applied due to the CRLB KPI becomes more complex when multiple targets or target-related bounced propagation paths need to be resolved. Moreover, besides of binary power loading within one OFDMA symbol, we can apply continues power loading [ref85]. This will be the case if only TDMA (hence no FDMA) multiplexing within the OFDMA frame is applied. While continuous power loading can be seen as equivalent to maximizing data transfer capacity respectively mutual information in multiuser OFDM (by the water filling algorithm), binary loading is equivalent to carrier avoidance power loading. The latter strategy prioritizes the use of high-performing subcarriers and disables low-performing subcarriers. The spared subcarriers can then be used for other links that can benefit more from the respective empty frequencies. For example, in a joint multilink measurement, the links with the highest resolution requirements in terms of delay and/or Doppler would be assigned REs in the outer regions of the frequency/time resource frame, while the links with lower resolution requirements would get REs more in the central part.
@@
The allocation of MS ISAC resources from the perspective of sensing alone already appears to be a complex optimization task. However, since the same radio resources are used for sensing and communication, the balance between communication and radar sensing performance must also be taken into account when scheduling resources. While in communication is all about the data rate and latency, in radar sensing the issue is how many dynamic targets can be reliably detected and tracked in a given scenario. Due to the dual use of radio resources inherent in the CPCL principle, the choice between communication and sensor technology regarding resource consumption is not a simple either/or since it is not even per se competitive. Moreover, cooperative sensing will entail an independent (additional) requirement for the transport of intermediate sensing data for fusion with special requirements in terms of latency and capacity. In addition, if there are not enough loadable carriers to provide required sensing performance, it would be appropriate to include dummy carriers that serve only for sensing. Eventually, MS ISAC resource scheduling would depend on the required sensing task and respective sensing cycle, making it a challenging joint multilink optimization task. The multi-criteria nature of the optimization problem yields a non-convex cost function that cannot be minimized by simple means. This suggests the need for an AI-supported solution. Another challenge results from the sparsity in frequency and time because of the multiple interleaved measurements. Therefore, estimation methods that can deal with sparse frames are required. A simple inverse filter and FFT application would fail. This is also due to the non-uniform carrier magnitudes due to multilevel modulation and power loading. The solution is model based target parameter estimation using weighted least squares as described above.
@@
As already described above, multipath propagation can be considered either a ’foe’, reducing target visibility and location accuracy, or a ’friend’, exploited to improve ISAC performance. Besides of explicit model based methods requiring a priori information from the geometric structure of the propagation environment and implicit in situ training based methods for multipath exploitation, there are frequency-time precoding schemes that use observed channel response functions and CSI feedback. These methods are known as Time Reversal Focusing (TRF) of waves. Originally, TRF was proposed by M. Fink in the context of ultrasonic waves [ref87]. Later on, it was also used to focus electromagnetic waves, with applications in communication and radar [ref88, ref89, ref34]. In short, the basic idea is as follows: When a spatial wave field is emitted from a source in a complex (’inhomogeneous’) medium and recorded by a set of sensors around the edge of the medium, then mirrored along the time axis (’time reversal’) and retransmitted, all wave components will refocus at the source, despite the time delay to the sensors and regardless of dispersive propagation effects (multipath reflections, scattering, diffraction, etc.). In communications, a basic application scenario for TRF assumes a UE that estimates the channel $H(f,t)$ as $\hat{H}(f,t)$ in the DL and transmits CSI to the gNB, where it is mirrored in time and convolved with the waveform to be transmitted. This process can be formulated in frequency and time domain as
@@
where $\hat{R_{H}}(f,t)$ denotes an estimate of the Fourier Transform of the channel’s Auto-Correlation Function (ACF) at time $t$ and $\hat{r_{h}}(\tau,t)$ its (fast) time domain counterpart.
@@
TR is always a two-step predictive procedure that relies on CSI estimation in DL and signaling in UL assuming that the channel does not change too fast. TR precoding changes the resulting channel impulse response to the ACF of the unmodified response (V-E).
@@
It appears that the focusing effect becomes better the stronger the delay spread is. Therefore, reverberant multipath rich channels transform to low delay spreads channels by precoding (as the channel ACF becomes short). In this sense, TCF can be interpreted as a “channel-matched filter”, which can help to simplify the receiver complexity by effectively shifting the equalization operation from the receiver to the transmitter. Obviously, focusing in delay goes along with true spatial focusing since the channel matched filter condition depends solely on the local superposition of multiple reflected paths impinging from different directions with different delays. This way, a multipath rich propagation environment acts like a huge reflect array (a TRF mirror) that causes a sharp spatial focus. If there is not enough multipath for focusing available, multiple cooperating MS ISAC transceivers or RIS can be deployed to achieve a similar effect. In addition to the shift in implementation effort from the UE to the infrastructure, there are further advantages, such as increased energy efficiency, since more reflected energy is collected and coherently combined at the Rx antenna. Moreover, in communications spatial focusing reduces the probability of non-desirable interception.
@@
Figure 20 shows an example of TR focusing calculated from measured channel impulse response functions. The sounding measurement took place in a NLOS street scenario in Cologne, Germany. The center frequency of the sounder was 3.75 GHz and bandwidth was 100 MHz. The sectorial BS-Antenna was above rooftop and the omnidirectional UE antenna was mounted on a slowly moving car on street level. The instantaneous power delay profiles (magnitude squared CIR) depicted in the upper left figure for two consecutive positions (green and orange) about 32 cm apart shows that LOS is obstructed. We can identify about 6…7 more or less relevant multipath components with one stronger reflection dominating. The green CIR was then taken as a TR reference depicted mirrored in delay in upper right figure. The resulting TR channel matched response is shown below (green). Note that in practice the response will appear in the right-hand side because of causality. Compared to the original CIR, we see a clear and strong focus in the delay domain, as was to be expected, since the original CIR has a noise-like structure. The orange response results from correlation of the same reference with the CIR at the displaced position. The degradation of about 10 dB in correlation gain because of the reference mismatch is obvious. This describes the expected spatial focusing. The CIR mismatch of the CIR is mainly because of the phase difference that is not visible in the power delay profile shown her. The distance of about 4 wavelengths leads to strong decorrelation, even if we only have this small number of paths with a somewhat limited angular spread.
@@
Figure 20: Time reversal focusing from measured CIRs. Upper row: Green/orange: Two measured CIRs, ca. 32 cm apart; Blue: Time reversed CIR from measured green CIR. Middle and lower row: TR channel matched response: Green matched to the correct CIR, orange matched to the displaced CIR.
@@
The situation in radar is somewhat different. Reason is that now we should focus on the passive target object in order to enhance its illumination and visibility against clutter. However, in most applications we will not have a transceiver for CSI sensing and reporting in place of the target. If we were now to perform end-to-end focusing of the transmit and receive antennas in Figure 6, we would obtain a focus at the receive antenna which, however, includes all multipath interference, that has nothing to do with the target. One promising solution to get rid of the not-target related clutter influence is background subtraction that removes or at least reduces the target independent clutter, including direct LOS. Background subtraction is easiest to use when the sensor network is static (e.g. in the case of infrastructure-only sensing) and the target is moving. It will become more complicated if the sensor nodes are moving and perhaps can be supported by beamforming. In any case, it is important that only propagation paths routed via the target are retained. If the end-to-end response is now precoded by time reversal, all and only the propagation paths that are routed via the target are used for the focusing at the sensor, see Figure 21.

### v2

@@ V-E Sensing Cycles – OFDMA Resource Allocation, Time-frequency Precoding and Link Adaptation
It seems beneficial to adjust ISAC radio access parameters and resources to maximize location and tracking parameters according to the current sensing mode. However, at the beginning of a reconnaissance mission, we generally have no a priori knowledge of the existence, position, and speed of the target. We would need dedicated radar sensing cycles following gradual levels of perception. In a first surveillance and acquisition cycle, the environment is scanned, and when a target is detected, quick, rough, and sometimes even incorrect estimates of the target parameters are made (e.g., many false alarms). Then, in a second step, the ISAC radio parameters will be adjusted to provide more accurate and less ambiguous estimates. This can be achieved, e.g., by a higher sensing symbol repetition rate, a longer coherent processing interval (c.f. radar integration time), which goes along with higher Doppler resolution, and by beamforming, which adapts to the directivity of propagation. However, since the dynamics and contributions of the multiple radio links depend on their relative positions with respect to the target, which change over time, ISAC-specific CSI estimation and feedback are needed for dynamic transmitter resource allocation, precoding, and link adaptation to support dynamic target tracking. Since this is a two-stage process based on previous CSI measurements, the effectiveness depends on the CSI pilot rate and the speed at which the channel changes.
@@
Link adaption in radar. Although the next generation mobile radio may include dedicated tools for ISAC quality of service control, it seems promising to explore how existing mechanisms and tools for CSI estimation and signaling can be used for ISAC performance control. Traditionally, intersymbol interference in mobile communications caused by multipath is compensated at the receiver by an equalizer. Later, it was realized that precoding at the transmitter side may be more effective, especially if multiple users are included. Over the decades of its maturation, mobile communications has developed sophisticated methods for precoding and resource allocation. Link adaptation has evolved into an unprecedentedly flexible multi-domain technique that includes scalable numerology, flexible resource element allocation in frequency and time, waveform selection, carrier power loading, modulation and coding, and multiple antenna precoding techniques including beamforming [ref47]. Together with advanced techniques for CSI estimation and signaling, the control channel allows reconfiguration and adaptation of the transmitter to the instantaneous link conditions for optimum quality of communication service. The latest stage has been set with 5G NR, which is designed to support enhanced mobile broadband, ultra-reliable low latency communications, and massive machine type communications. Since coordination of multiple links and multipath exploitation will be even more important for performance optimization in ISAC (due to the geometric relevance of the multiple distributed links), we expect that precoding and link adaptation will be of similar or even greater importance.
@@
Let’s take another quick look at the link adaptation in radar. The situation is slightly different here. Although multistatic and networked radars have a long-lasting history [ref49], single-station monostatic radars are still the default design [ref48]. The reasons for this seem to be manifold. Monostatic radars are independent and can act autonomously (even mobile). The underlying geometry, a circle or a sphere, seems to be simple. Most of all, signal processing is local, which makes Tx / Rx synchronization and resource adaptation (bandwidth, Tx power, pulse rate, dwell time, beam management, etc.) a local task. There are very advanced resource management principles for monostatic radars applied that can cope with dynamic scenarios of multiple targets [ref46] and are able to adapt their operation based on cognitive procedures [ref44]. Attempts to overcome the geometric limitations of coverage and target visibility by a network of radar sensors often end up with non-coherent cooperation of multiple monostatic radars [ref45] (fusion of multiple directions from dislocated radars is a non-coherent operation). References to radar networks with multi-static cooperation are rare [ref48]. Even if the publicly available literature may not completely reflect the entire state of knowledge, the radar-centric approach to ISAC, which requires a separate (probably proprietary) communication network or additional communication functions integrated into the radar’s air interface, does not appear to be a very promising role model for large scale application of ISAC in public networks. At the very least, ISAC’s communications-centric approach seems more promising, as it reuses the already available communication network for data transport and fusion. In addition, the well-designed and constantly evolving sensing, signaling, and control schemes of mobile communications reveal a kind of hitchhiking support for radar functionality. This is most obvious for the CPCL approach, which seems to be the closest coexistence of communication and sensing.
@@
Radio access coordination of multiple sensing links. As we have explained above, ISAC radar detection and tracking of 3D dynamic target state vectors in general requires multiple sensing links that have to be coordinated and adapted to the dynamic target scenario to meet the required sensing performance. It is obvious that resource management and link adaptation must take place in a distributed MS-ISAC system at the radio and network level. While the primary goal of communication is efficient transfer of data between two dedicated radio nodes acting as the source and the destination, the goal of ISAC is the localization and recognition of target objects. Multiuser communication links, in essence, can be seen as independent single links just sharing common radio resources according to rules of fairness or priority. Sensing rather requires the proactive cooperation of multiple radio nodes (illuminators and sensors) that share the same resources and work together to fulfill a common sensing task. Already when selecting the nodes involved, their spatial availability (both current and predicted), expected influence on localization accuracy due to GDoP and SNR, and detection functionality due to heterogeneous configuration and connectivity have to be taken into account. Furthermore, the allocation of communication and computation resources is based on the decision on the data collection task and the type of data fusion. It seems important to recognize that not a single radio node (e.g., a UE) is a beneficiary of an ISAC service, but the fused information from several sensing links that is made available to authorized users represents the added value. Therefore, MS-ISAC resource allocation is always a joint, respectively cooperative activity of a sensing network.
@@
Once we have defined which radio nodes contribute to the execution of a sensing mission, we need to decide how to allocate radio access resources. This is especially interesting when multiple sensing links share the same radio channel. Which signals and access schemes do we have available for target scene illumination in ISAC? While dedicated pilot signals are required to establish communication links (Tx / Rx synchronization, demodulation reference signaling, and CSI estimation), and therefore are essential also for ISAC operation, the use of CSI pilots alone will in general not be enough for ISAC sensing as already explained in the subsection on Recovery of the Correlation Reference Signal. The use of the Positioning Reference Signal would be another option. The PRS is a dedicated signal that enables self-localization of mobile UEs in the DL. It can carry network-specific data to support localization services and allows OFDMA frame mapping to support multiple UEs [ref74]. Although it is not intended for radar sensing, it can be used in the DL of MS-ISAC to temporarily enhance target illumination. However, excessive use of PRS would only block resources that may not be available for communication services. This means that radio resource efficient ISAC should aim to reuse the communication payload for sensing, as already proposed by the CPCL principle. It is also possible to use the communication payload-based sensing results for CSI feedback and sensing link adaptation. The use of CSI pilots has the advantage that orthogonal multi-sensor access is easy to achieve.
@@
Broadcast and multi-sensor access. According to the subsection on Multi-Sensor Radio Access and Scheduling, it is reasonable to distinguish between the broadcast mode and multi-sensor ISAC access. In the first case, each transmitter serves several sensing receivers, which provides simultaneous measurements and, thus, instantaneously multiplies the information about the target. This applies, e.g., if the sensing nodes are receive-only sniffers deployed as part of a distributed infrastructure or acting as deployable sniffers in the reduced CPCL-mode as described in section II. In case of MS-ISAC, the orthogonal multiplexed sensing links must share the same channel by proper allocation of the RBs in the OFDM frequency-time frame as described in the subsection on Multi-Sensor Radio Access and Scheduling. Obviously, resource sharing should be made adaptive according to the number and dynamics of targets, as well as the multipath propagation situation.
@@
For the purpose of MS-ISAC radio access in DL or UL, we must adjust the distribution of RBs in the frequency-time frame by OFDMA. From the instantaneous ISAC CSI and KPI estimates transmitted via the UL to the Central Unit (CU) of a distributed base station, a resource scheduler can decide how many resources per link have to be spent to follow the dynamically changing parameters of the respective target. Later on, when the moving target is tracked, and the dynamic situation changes according to the target’s position and velocity vector relative to the baseline of the sensor links, the RB distribution will be updated.
@@
CRB driven allocation of OFDMA resource blocks. Since the target resolution performance depends on the CRB as discussed in the subsection on Model-based Bistatic Target Parameter Estimation, it is recommended to use it as a cost function for MS-ISAC resource scheduling. For the MS access, the bottom line is about optimum arrangement and average power weighting of carriers in the OFDMA frame for any sensing link. The total number of carriers to be allocated per link depends on both the required processing gain of the matched filter (according to the SNR of the respective link) and the resolution performance. The CRB criterion tends to shift the carriers in blocks to the edges of the available frequency-time frame. Since we may have to host multiple sensing links, we would need to interleave them in two dimensions, frequency and time. A simple rule of thumb for RB scheduling of multiple cooperating sensing links may be to allocate the RBs of links with temporarily lower resolution requirements to the inner space of the OFDMA frame. The advantage of the 2D nesting is that virtual simultaneous measurements from multiple transmitters become possible, with the restrictions that a shared medium entails.
@@
The weighting applied due to the CRB KPI becomes more complex when multiple targets or target-related bounced propagation paths need to be resolved. Moreover, besides binary power loading within one OFDMA symbol, we can apply continuous power loading [ref85]. This will be the case if only TDMA (hence no FDMA) multiplexing within the OFDMA frame is applied. While continuous power loading can be seen as equivalent to maximizing data transfer capacity, respectively mutual information, in multiuser OFDM (by the water filling algorithm), binary loading is equivalent to carrier avoidance power loading. The latter strategy prioritizes the use of high-performing subcarriers and disables low-performing subcarriers. The spared subcarriers can then be used for other links that can benefit more from the respective empty frequencies. For example, in a joint multilink measurement, the links with the highest resolution requirements in terms of delay and / or Doppler would be assigned REs in the outer regions of the frequency/time resource frame, while the links with lower resolution requirements would get REs more in the central part.
@@
AI enabled resource allocation. The allocation of MS-ISAC resources from the perspective of sensing alone already appears to be a complex optimization task. However, since the same radio resources are used for sensing and communication, the balance between communication and radar sensing performance must also be taken into account when scheduling resources. While in communication, it is all about the data rate and latency; in radar sensing, the issue is how many dynamic targets can be reliably detected and tracked in a given scenario. Due to the dual use of radio resources inherent in the CPCL principle, the choice between communication and sensor technology regarding resource consumption is not a simple either / or since it is not even per se competitive. Moreover, cooperative sensing will entail an independent (additional) requirement for the transport of intermediate sensing data for fusion with special requirements in terms of latency and capacity. In addition, if there are not enough loadable carriers to provide the required sensing performance, it would be appropriate to include dummy carriers that serve only for sensing. Eventually, MS-ISAC resource scheduling would depend on the required sensing task and respective sensing cycle, making it a challenging joint multilink optimization task. The multi-criteria nature of the optimization problem yields a non-convex cost function that cannot be minimized by simple means. This suggests the need for an AI-supported solution. Another challenge results from the sparsity in frequency and time because of the multiple interleaved measurements. Therefore, estimation methods that can deal with sparse frames are required. A simple inverse filter and FFT application would fail. This is also due to the non-uniform carrier magnitudes due to multilevel modulation and power loading. The solution is model-based target parameter estimation using weighted least squares as described above.
@@
Multipath exploiting time reversal precoding. As already described above, multipath propagation can be considered either a ’foe’, reducing target visibility and location accuracy, or a ’friend’, exploited to improve ISAC performance. Besides explicit model-based methods requiring a priori information from the geometric structure of the propagation environment and implicit in situ training-based methods for multipath exploitation, there are frequency-time precoding schemes that use observed channel response functions and CSI feedback. These methods are known as Time Reversal Focusing (TRF) of waves. Originally, TRF was proposed by M. Fink in the context of ultrasonic waves [ref87]. Later on, it was also used to focus electromagnetic waves, with applications in communication and radar [ref88, ref89, ref34]. In short, the basic idea is as follows: When a spatial wave field is emitted from a source in a complex (’inhomogeneous’) medium and recorded by a set of sensors around the edge of the medium, then mirrored along the time axis (’time reversal’) and retransmitted, all wave components will refocus at the source, despite the time delay to the sensors and regardless of dispersive propagation effects (multipath reflections, scattering, diffraction, etc.). In communications, a basic application scenario for TRF assumes a UE that estimates the channel $H(f,t)$ as $\hat{H}(f,t)$ in the DL and transmits the CSI to the gNB, where it is mirrored in time and convolved with the waveform to be transmitted. This process can be formulated in the frequency and time domains as
@@
where $\hat{R_{H}}(f,t)$ denotes an estimate of the FT of the channel’s Auto-Correlation Function (ACF) at time $t$ and $\hat{r_{h}}(\tau,t)$ its (fast) time domain counterpart.
@@
Time Reversal (TR) is always a two-step predictive procedure that relies on the CSI estimation in DL and signaling in UL, assuming that the channel does not change too fast. TR precoding changes the resulting channel impulse response to the ACF of the unmodified response (V-E).
@@
It appears that the focusing effect becomes better the stronger the delay spread is. Therefore, reverberant multipath rich channels transform to low delay spread channels by precoding (as the channel ACF becomes short). In this sense, TRF can be interpreted as a “channel-matched filter”, which can help to simplify the receiver complexity by effectively shifting the equalization operation from the receiver to the transmitter. Obviously, focusing in delay goes along with true spatial focusing since the channel-matched filter condition depends solely on the local superposition of multiple reflected paths impinging from different directions with different delays. This way, a multipath-rich propagation environment acts like a huge reflect array (a TRF mirror) that causes a sharp spatial focus. If there is not enough multipath for focusing available, multiple cooperating MS-ISAC transceivers or RIS can be deployed to achieve a similar effect. In addition to the shift in implementation effort from the UE to the infrastructure, there are further advantages, such as increased energy efficiency, since more reflected energy is collected and coherently combined at the Rx antenna. Moreover, in communications, spatial focusing reduces the probability of non-desirable interception.
@@
Figure 20 shows an example of TR focusing calculated from measured channel impulse response functions. The sounding measurement took place in a Non Line of Sight (NLOS) street scenario in Cologne, Germany. The center frequency of the sounder was $3.75\text{\,}\mathrm{GHz}\text{/}$ and the bandwidth was $100\text{\,}\mathrm{MHz}\text{/}$. The sectorial BS-Antenna was above the rooftop, and the omnidirectional UE antenna was mounted on a slowly moving car on street level. The instantaneous power delay profiles (magnitude squared CIR) depicted in the upper left figure for two consecutive positions (green and orange) about 32 cm apart show that LOS is obstructed. We can identify about 6 to 7 more or less relevant multipath components, with one strong reflection dominating. The green CIR was then taken as a TR reference depicted mirrored in delay in the upper right figure. The resulting TR channel matched response is shown below (green). Note that in practice, the response will appear on the right-hand side because of causality. Compared to the original CIR, we see a clear and strong focus in the delay domain, as was to be expected, since the original CIR has a noise-like structure. The orange response results from the correlation of the same reference with the CIR at the displaced position. The degradation of about 10 dB in correlation gain because of the reference mismatch is obvious. This describes the expected spatial focusing. The CIR mismatch of the CIR is mainly because of the phase difference that is not visible in the power delay profile shown here. The distance of about 4 wavelengths leads to strong decorrelation, even if we only have this small number of paths with a somewhat limited angular spread.
@@
Figure 20: Time reversal focusing from measured CIRs. Upper row: Green/orange: Two measured CIRs, ca. $32\text{\,}\mathrm{cm}\text{/}$ apart; Blue: Time reversed CIR from measured green CIR. Middle and lower row: TR channel-matched response: Green matched to the correct CIR, orange matched to the displaced CIR.
@@
The situation for radar is somewhat different. The reason is that now we should focus on the passive target object in order to enhance its illumination and visibility against clutter. However, in most applications, we will not have a transceiver for CSI sensing and reporting in place of the target. If we were now to perform end-to-end focusing of the transmit and receive antennas in Figure 6, we would obtain a focus at the receive antenna, which, however, includes all multipath interference, that has nothing to do with the target. One promising solution to get rid of the not-target-related clutter influence is background subtraction that removes or at least reduces the target-independent clutter, including direct LOS. Background subtraction is easiest to use when the sensor network is static (e.g., in the case of infrastructure-only sensing) and the target is moving. It will become more complicated if the sensor nodes are moving, and perhaps can be supported by beamforming. In any case, only propagation paths routed via the target must be retained. If the end-to-end response is now precoded by time reversal, all and only the propagation paths that are routed via the target are used for the focusing at the sensor, see Figure 21.